# Probabilistic Approaches to The Energy Equality in Forced Surface Quasi-Geostrophic Equations

### v1

@@ Abstract.
We explore probabilistic approaches to the deterministic energy equality for the forced Surface Quasi-Geostrophic (SQG) equation on a torus. First, we prove the zero-noise dynamical large deviations for a corresponding stochastic SQG equation, where the lower bound matches the upper bound on a certain closure of the weak-strong uniqueness class for the deterministic forced SQG equation. Furthermore, we show that the energy equality for the deterministic SQG equation holds on arbitrary time-reversible subsets of the domain, where we match the upper bound and the lower bound. Conversely, the violation of the deterministic energy equality breaks the lower bound of large deviations. These results extend the existing techniques in Gess, Heydecker, and the second author [GHW24] to generalized Sobolev spaces with negative indices. Finally, we provide an analysis of the restricted quasi-potential and prove a conditional equivalence compared to the rate function of large deviations for the Gaussian distribution. This suggests a potential connection between non-Gaussian large deviations in equilibrium for the stochastic SQG equation and the open problem regarding the uniqueness of the deterministic SQG equation.
@@
Surface Quasi-Geostrophic equation, large deviations, energy equality.
@@ 1. Introduction
within the framework of large deviations theory. Here, $\xi_{\delta(\varepsilon)}$ is a spatial regularization of space-time white noise $\xi$. We will explore the connections between dynamical large deviations of (1.3) as $(\varepsilon,\delta(\varepsilon))\rightarrow(0,0)$ and the validity of the deterministic energy equality in (1.1). Additionally, we will highlight the potential relationships between large deviations in equilibrium for (1.3) and the uniqueness problem for (1.1). Furthermore, when $\varepsilon=0$ in (1.3), a scaling argument suggests that the fractional dissipation index $\alpha=1/2$ belongs to the critical regime, $\alpha>1/2$ falls within the subcritical regime and $\alpha\in(0,1/2)$ stands for the supercritical regime [Res95]. Hence the ranges in (1.2) cover subcritical, critical, and supercritical cases.
@@
holds for a Leray solution of (1.1). Analogous to the celebrated result of Lions-Ladyzhenskaya for three-dimensional Navier-Stokes equations, it is possible to verify the energy equality for the SQG equation under some regularity conditions, see for example [Dai17]. On the other hand, the uniqueness problem of (1.1) remains open as well, with various regularity conditions specified to show uniqueness. We refer readers to [Res95, Section 3.4], [DC06] and the references therein for more details. To the best of our knowledge, in the study of SQG equation, there are no direct relationships between the regularity conditions for the energy equality and those for uniqueness. This is one of the motivations for this work, and we point out that a direct relationship will be provided in this context based on the perspective of probability. More precisely, this is linked to the study of dynamical large deviations of (1.3). Furthermore, this unveils a sufficient condition for proving the energy equality (1.4), although it is unclear how to verify such a condition.
@@
The stochastic PDE (1.3) (proposed by Totz [Tot20] and Hofmanová et al. [HLZZ24]) is an SQG equation driven by additive noise with small intensity $\varepsilon$ and small correlation $\delta(\varepsilon)$. The consideration of such noise follows from the principles of fluctuating hydrodynamics theory. Generally speaking, the SQG equation can be regarded as a toy model for understanding the regularity of the Navier-Stokes equations. We point out that an important fluctuating Navier-Stokes model known as the Landau-Lifshitz-Navier-Stokes (or Navier-Stokes-Fourier system for the full equations) is governed by the fluctuation-dissipation principle, a fundamental concept in the theory of fluctuating hydrodynamics (see [Ö05]). The stochastic SQG equation (1.3) could be considered as a stochastic toy model for the Landau-Lifshitz-Navier-Stokes equations. In particular, the choice of $\beta=\frac{\alpha}{2}$, corresponding to the fractional dissipation $-\Lambda^{2\alpha}$, exhibits a structural resemblance to the fluctuation-dissipation principle. Regarding (1.3), a dynamical large deviation principle is proved, in a joint scaling regime $(\varepsilon,\delta(\varepsilon))\rightarrow(0,0)$, with initial data allowing for fluctuations as well. Moreover, we provide an analysis on the identification of the quasi-potential
@@
$\mathcal{U}(\phi)=\frac{1}{2}\inf\Big{\{}\int^{0}_{-\infty}\|\partial_{t}\tilde{\theta}+\Lambda^{2\alpha}\tilde{\theta}+R^{\perp}\tilde{\theta}\cdot\nabla\tilde{\theta}\|_{H^{-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s:\ -\tilde{\theta}(-\cdot)\in\mathcal{A}_{0},\ \tilde{\theta}(0)=\phi\Big{\}},$ (1.5)
@@
Here, we emphasize that in the study of large deviations for (1.3), the initial data could also be considered in $\dot{H}^{-1/2}(\mathbb{T}^{2})$. In such a lower regularity regime, a commutator estimate associated with the Riesz transform is employed to make the nonlinear term well-defined. This distinguishes the technique used in proving large deviations for SQG equations from that for three-dimensional Landau-Lifshitz-Navier-Stokes equations established by Gess, Heydecker, and the second author [GHW24]. Further comments regarding the technique and the physical background for such an $\dot{H}^{-1/2}(\mathbb{T}^{2})$ framework will be provided later on.
@@
(UL): Weak-strong uniqueness implies a restricted dynamical large deviation principle;
(LE): The energy equality holds on any time-reversible domain where the lower bound of large deviations matches the upper bound;
(EL): Violation of the kinetic energy equality breaks the lower bound of large deviations;
(EU): The energy equality holds on a certain time-reversible closure of the weak-strong uniqueness regularity class;
(UN): The conditional equivalence of the quasi-potential and the Gaussian rate function is proved.
We explain why (UN) provides a potential connection between non-Gaussian large deviations in equilibrium and open problems of uniqueness for deterministic SQG equation. In general, the proof of a large deviation principle for invariant measures of stochastic PDEs relies on uniform dynamical large deviations, and the rate function is governed by the quasi-potential, which is defined as the infimum of the dynamical cost on the time interval $(-\infty,0]$ along paths in the space where the dynamical large deviations are proved. However, in this paper, the dynamical large deviation lower bound is shown to match the upper bound only when we restrict fluctuations to a closure of the so-called weak-strong uniqueness class $\mathcal{C}_{0}$ (see Definition 1.1 later on). This suggests that the quasi-potential should be defined in a restricted version, see (1.5), where the infimum is taken along paths corresponding to a subset $\mathcal{A}_{0}\subset\mathcal{C}_{0}$. To show that the quasi-potential (1.5) equals the Gaussian rate function $I_{Gauss}(\phi)=\|\phi\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$ for all $\phi\in H^{\alpha-2\beta}(\mathbb{T}^{2})$, a similar argument as in [BCF15] can be carried out. However, this relies on the existence of solutions for the deterministic SQG equation in $\mathcal{A}_{0}$. This links to the open problem of the uniqueness for the SQG equation.
@@ 1.1. Main results
where ”$w$” denotes the weak topology. Let $s>\alpha+1$, we regularize the noise by
$\xi_{\delta}=\sqrt{Q_{\delta}}\xi=(I+\delta\Lambda^{2s})^{-1/2}\xi.$
A detailed calculation concerning the blow-up speed of $\xi_{\delta}$ will be given in Section 2. For every $\theta\in\mathbb{X}_{\alpha,\beta}$, let
$\mathcal{I}_{dyna}(\theta)=\frac{1}{2}\bigg{\|}\partial_{t}\theta+\Lambda^{2\alpha}\theta+R^{\perp}\theta\cdot\nabla\theta\bigg{\|}^{2}_{L^{2}\left([0,T];H^{-2\beta}(\mathbb{T}^{2})\right)},\ \ \ \mathcal{I}_{0}(\theta(0))=\|\theta(0)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$
@@
We say that $\mathcal{C}_{0}\subset\mathbb{X}_{\alpha,\beta}$ is a weak-strong uniqueness class of (1.1) for every $g\in L^{2}([0,T];L^{2}(\mathbb{T}^{2}))$ if, for arbitrary two weak solutions $\theta_{1},\theta_{2}$ (see Definition 4.1 later on) of the skeleton equation (1.1) with the same initial data $\theta_{1}(0)=\theta_{2}(0)$ and the same control $g$, we have $\theta_{1}=\theta_{2}$ in $\mathbb{X}_{\alpha,\beta}$ as long as $\theta_{1}\in\mathcal{C}_{0}$ and $\theta_{2}$ satisfies the $H^{\alpha-2\beta}$-energy inequality: for every $t\in[0,T],$
@@
$\overline{\mathcal{C}_{0}}^{\mathcal{I}}:=\big{\{}\theta\in\mathbb{X}_{\alpha,\beta}:\text{there exists a sequence }\{\theta^{(n)}\}\subset\mathcal{C}_{0},\,\text{such that }\theta^{(n)}\rightarrow\theta\text{ in }\mathbb{X}_{\alpha,\beta},\,\mathcal{I}\big{(}\theta^{(n)}\big{)}\rightarrow\mathcal{I}(\theta)\big{\}}.$
@@
Assume that the law of the initial data $\theta_{\varepsilon,\delta}(0)$ is $\mathcal{G}\left(0,\varepsilon Q_{\delta}/2\right)$, the Gaussian measure on $\dot{H}^{\alpha-2\beta}(\mathbb{T}^{2})$ with mean zero and covariance $\varepsilon Q_{\delta}/2.$
Remark 1.2.
Compared to the classical large deviations for stochastic PDEs, we allow for fluctuations in the initial data, which constitutes a theoretical generalization. In particular, for the case $\delta=0$, solutions of the Galerkin approximation equation of (1.3) starting from equilibrium exhibit the time-reversibility (see Lemma 7.2 below). We will later see how time-reversibility is related to the kinetic energy equality. We further remark that the above assumption can be extended to more general conditions (see [GHW24, Assumption 2.1]).
@@
For every $\varepsilon,\delta>0$, let $\theta_{\varepsilon,\delta}$ be a stochastic generalized Leray solution of (1.3) in the sense of Definition 3.1 with initial data $\theta_{\varepsilon,\delta}(0)$ satisfying Assumption A. Let $\mu_{\varepsilon}=\mu_{\varepsilon,\delta}$ be the laws of $\theta_{\varepsilon,\delta}$ on $\mathbb{X}_{\alpha,\beta}$. Let $\mathcal{C}_{0}\subset\mathbb{X}_{\alpha,\beta}$ be a weak-strong uniqueness class of (1.1) and $\mathcal{C}=\overline{\mathcal{C}_{0}}^{\mathcal{I}}$. Assume that the scaling regime (2.7) holds for $(\varepsilon,\delta(\varepsilon))$. Then for any closed set $F\subset\mathbb{X}_{\alpha,\beta}$,
@@
We further comment that the above large deviations result holds for a sequence of Galerkin approximations of (1.3) (See (3.7) below) as well, see Proposition 5.1 and Proposition 6.3 for details. Based on this large deviations result, we provide a direct relationship between regularity classes of the weak-strong uniqueness and regularity classes where the energy equality holds. Concretely, we show that the kinetic energy equality holds on every time-reversible set of a certain closure of the weak-strong uniqueness class $\mathcal{C}_{0}$. This introduces the following result.
@@
Assume that $\beta=\alpha/2$. Let $\mathcal{C}_{0}$ be a weak-strong uniqueness class of (1.1) and $\mathcal{C}=\overline{\mathcal{C}_{0}}^{\mathcal{I}}$. Let $\mathcal{R}\subset\mathcal{C}$ satisfy $\mathcal{R}=\mathfrak{T}_{T}\mathcal{R}$, where the time-reversal operator $\mathfrak{T}_{T}:\mathbb{X}_{\alpha,\alpha/2}\rightarrow\mathbb{X}_{\alpha,\alpha/2}$ is defined by $(\mathfrak{T}_{T}\theta)(\cdot):=-\theta(T-\cdot)$. Suppose that $\theta\in\mathcal{R}$ is a weak solution of (1.1) for some $g\in L^{2}([0,T];L^{2}(\mathbb{T}^{2}))$ in the sense of Definition 4.1, then the following kinetic energy equality holds:
@@
Let $\varepsilon>0$ and $m\in\mathbb{N}_{+}$ satisfy the scaling regime (2.8). Let $\theta_{\varepsilon,m}$ be a solution of (1.10) with initial data $\theta_{\varepsilon,m}(0)\sim\mathcal{G}(0,\varepsilon P_{m}/2)$ and let $\mu_{\varepsilon,m}$ be the law of $\theta_{\varepsilon,m}$ on $\mathbb{X}_{\alpha,\alpha/2}.$ Suppose that $\theta$ is a weak solution of (1.1) for some $g\in L^{2}([0,T];L^{2}(\mathbb{T}^{2}))$ in the sense of Definition 4.1.
@@
$\inf_{\{G\text{ is open in }\mathbb{X}_{\alpha,\alpha/2}:\theta\in G\}}\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon,m}(G)<-\mathcal{I}(\theta).$
@@
$\inf_{\{G\text{ is open in }\mathbb{X}_{\alpha,\alpha/2},\,\mathfrak{T}_{T}\theta\in G\}}\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon,m}(G)<-\mathcal{I}(\mathfrak{T}_{T}\theta).$
Finally, we provide an analysis on the exact representation of the quasi-potential (1.5). Let $\mathcal{C}_{0}$ be a weak-strong uniqueness class on $[0,\infty)$ (See Definition 8.2). Denote
$\mathcal{A}_{0}:=\big{\{}\theta\in\mathcal{C}_{0}:\partial_{t}\theta\in L^{2}([0,\infty);H^{\alpha-2\beta}(\mathbb{T}^{2})),R^{\perp}\theta\cdot\nabla\theta\in L^{2}([0,\infty);H^{\alpha-2\beta}(\mathbb{T}^{2}))$ ,
[…]

### v3

@@ Abstract.
We explore probabilistic approaches to the deterministic energy equality for the forced Surface Quasi-Geostrophic (SQG) equation on a torus. First, we prove the zero-noise dynamical large deviations for a corresponding stochastic SQG equation, where the lower bound matches the upper bound on a certain closure of the weak-strong uniqueness class for the deterministic forced SQG equation. Furthermore, we show that the energy equality for the deterministic SQG equation holds on arbitrary time-reversible subsets of the domain where we match the upper bound and the lower bound. Conversely, the violation of the deterministic energy equality breaks the lower bound of large deviations. These results extend the existing techniques in Gess, Heydecker, and the second author [GHW24_LLNS] to generalized Sobolev spaces with negative indices. Finally, we provide an analysis of the restricted quasi-potential and prove a conditional equivalence compared to the rate function of large deviations for the Gaussian distribution. This suggests a potential connection between non-Gaussian large deviations in equilibrium for the stochastic SQG equation and the open problem regarding the uniqueness of the deterministic SQG equation.
@@
Surface Quasi-Geostrophic equation, large deviations, energy equality, weak-strong uniqueness class, non-Gaussian equilibrium.
@@ 1. Introduction
within the framework of large deviations theory. Here, $\xi_{\delta(\varepsilon)}$ is a spatial regularization of space-time white noise $\xi$. We will explore the connections between dynamical large deviations of (1.3) as $(\varepsilon,\delta(\varepsilon))\rightarrow(0,0)$ and the validity of the deterministic energy equality in (1.1). Additionally, we will highlight the potential relationships between large deviations in equilibrium for (1.3) and the uniqueness problem for (1.1). Furthermore, when $\varepsilon=0$ in (1.3), a scaling argument suggests that the fractional dissipation index $\alpha=1/2$ belongs to the critical regime, $\alpha>1/2$ falls within the subcritical regime, and $\alpha\in(0,1/2)$ stands for the supercritical regime [Res95]. Hence the ranges in (1.2) cover subcritical, critical, and supercritical cases.
@@
holds for a Leray solution of (1.1). In the system of the SQG equation, the field $\theta$ represents the temperature or surface buoyancy for a rapidly rotating stratified fluid with uniform potential vorticity, and $u_{\theta}=R^{\perp}\theta$ denotes the transport velocity field [HPGS95]. From a mathematical perspective, the energy equality (1.4) is expected to hold formally because $u_{\theta}$ is divergence-free. For sufficiently regular $\theta$, taking the $L^{2}$-inner product of both sides of equation (1.1) with $\theta$ then yields (1.4). Moreover, since the Riesz transform is an isometry on $L^{2}$ [SW71, Chapter VI, Section 2], the $L^{2}$-norm of $\theta$ equals that of $u_{\theta}$ and thus corresponds to the kinetic energy
@@
Analogous to the celebrated result of Lions-Ladyzhenskaya for three-dimensional Navier-Stokes equations, it is possible to verify the energy equality for the SQG equation under some regularity conditions, see for example [Dai17_energy]. On the other hand, the uniqueness problem of (1.1) remains open as well, with various regularity conditions specified to show uniqueness. We refer readers to [Res95, Section 3.4], [DC06_weak_strong], and the references therein for more details. To the best of our knowledge, in the study of SQG equation, there are no direct relationships between the regularity conditions for the energy equality and those for uniqueness. This is one of the motivations for this work, and we point out that a direct relationship will be provided in this context based on the perspective of probability. More precisely, this is linked to the study of dynamical large deviations of (1.3). Furthermore, this unveils a sufficient condition for proving the energy equality (1.4), although it is unclear how to verify such a condition.
@@
The stochastic PDE (1.3) (proposed by Totz [Tot20] and Hofmanová et al. [HLZZ24]) is an SQG equation driven by additive noise with small intensity $\varepsilon$ and small correlation $\delta(\varepsilon)$. The consideration of such noise is inspired by the fluctuation-dissipation relation at a formal level. Generally speaking, the SQG equation can be regarded as a toy model for understanding the regularity of the Navier-Stokes equations. We point out that an important fluctuating Navier-Stokes model known as the Landau-Lifshitz-Navier-Stokes (or Navier-Stokes-Fourier system for the full equations) is governed by the fluctuation-dissipation principle, a fundamental concept in the theory of fluctuating hydrodynamics (see [O05]). The stochastic SQG equation (1.3) could be considered as a stochastic toy model for the Landau-Lifshitz-Navier-Stokes equations. In particular, the choice of $\beta=\frac{\alpha}{2}$, corresponding to the fractional dissipation $\Lambda^{2\alpha}$, exhibits a structural resemblance to the fluctuation-dissipation principle. Regarding (1.3), a dynamical large deviation principle is proved, in a joint scaling regime $(\varepsilon,\delta(\varepsilon))\rightarrow(0,0)$, with initial data allowing for fluctuations as well. Moreover, we provide an analysis on the identification of the quasi-potential
@@
$\mathcal{U}(\phi)=\frac{1}{2}\inf\bigg\{\int^{0}_{-\infty}\big\|\partial_{t}\tilde{\theta}+\Lambda^{2\alpha}\tilde{\theta}+R^{\perp}\tilde{\theta}\cdot\nabla\tilde{\theta}\big\|_{H^{-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s:\ -\tilde{\theta}(-\cdot)\in\mathcal{A}_{0},\ \tilde{\theta}(0)=\phi\bigg\},$ (1.5)
@@
Here, we emphasize that in the study of large deviations for (1.3), the initial data could also be considered in $\dot{H}^{-1/2}(\mathbb{T}^{2})$. For the inviscid SQG equation, the functional
can be taken as the Hamiltonian of the system. This is derived from the Euler-Poincaré variational principle. We refer readers to Arnold’s work [Arn66], which is an infinite-dimensional generalization of [Poi01]. (See [Res95, Section 2.2], [BSV19_Nonuniqueness_SQG, Section 1.4, Appendix A.1] for more details.) In such a lower regularity regime, a commutator estimate associated with the Riesz transform is employed to make the nonlinear term well-defined. This distinguishes the technique used in proving large deviations for SQG equations from that for three-dimensional Landau-Lifshitz-Navier-Stokes equations established by Gess, Heydecker, and the second author [GHW24_LLNS]. Further comments regarding the technique for such an $\dot{H}^{-1/2}(\mathbb{T}^{2})$ framework will be provided later on.
@@
(UL): (Theorem 1.3) Weak-strong uniqueness implies a restricted dynamical large deviation principle;
(LE): (Theorem 1.4, probabilistic approach) The energy equality holds on any time-reversible domain where the lower bound of large deviations matches the upper bound;
(EL): (Corollary 1.5) Violation of the kinetic energy equality breaks the lower bound of large deviations;
(EU): (Theorem 1.4, Theorem 7.1, analytic approach) The energy equality holds on a certain time-reversible closure of the weak-strong uniqueness regularity class;
(UN): (Theorem 1.6) The conditional equivalence of the quasi-potential and the Gaussian rate function is proved.
We explain why (UN) provides a potential connection between non-Gaussian large deviations in equilibrium and the open problems of uniqueness for the deterministic SQG equation. In general, the proof of a large deviation principle for invariant measures of stochastic PDEs relies on uniform dynamical large deviations, and the rate function is governed by the quasi-potential, which is defined as the infimum of the dynamical cost on the time interval $(-\infty,0]$ along paths in the space where the dynamical large deviations are proved. However, in this paper, the dynamical large deviation lower bound is shown to match the upper bound only when we restrict fluctuations to a closure of the so-called weak-strong uniqueness class $\mathcal{C}_{0}$ (see Definition 1.1 later on). This suggests that the quasi-potential should be defined in a restricted version, see (1.5), where the infimum is taken along paths corresponding to a subset $\mathcal{A}_{0}\subset\mathcal{C}_{0}$. To show that the quasi-potential (1.5) equals the Gaussian rate function $I_{Gauss}(\phi)=\|\phi\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$ for all $\phi\in H^{\alpha-2\beta}(\mathbb{T}^{2})$, a similar argument as in [BCF15] can be carried out. However, this relies on the existence of solutions for the deterministic SQG equation in $\mathcal{A}_{0}$. This links to the open problem of the uniqueness for the SQG equation.
@@ 1.1. Main results
where “ww” denotes the weak topology. Let $s>\alpha+1$, we regularize the noise by
$\xi_{\delta(\varepsilon)}=\sqrt{Q_{\delta(\varepsilon)}}\xi=(I+\delta(\varepsilon)\Lambda^{2s})^{-1/2}\xi.$
A detailed calculation concerning the blow-up speed of $\xi_{\delta(\varepsilon)}$ will be given in Section 2. Based on this, we restrict that the scaling regime $(\varepsilon,\delta(\varepsilon))$ satisfies
$\lim_{\varepsilon\rightarrow 0}\varepsilon\delta(\varepsilon)^{-\frac{\alpha+1}{s}}=0$ (1.7)
in the study of large deviations. For every $\theta\in\mathbb{X}_{\alpha,\beta}$, let
$\mathcal{I}_{dyna}(\theta)=\frac{1}{2}\big\|\partial_{t}\theta+\Lambda^{2\alpha}\theta+R^{\perp}\theta\cdot\nabla\theta\big\|^{2}_{L^{2}\left([0,T];H^{-2\beta}(\mathbb{T}^{2})\right)},\ \ \ \mathcal{I}_{0}(\theta(0))=\|\theta(0)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$
@@
We say that $\mathcal{C}_{0}\subset\mathbb{X}_{\alpha,\beta}$ is a weak-strong uniqueness class of (1.1) if for every control $g\in L^{2}([0,T];L^{2}(\mathbb{T}^{2}))$, the following holds: for arbitrary two weak solutions (see Definition 4.1 later on) $\theta_{1},\theta_{2}$ of the skeleton equation (1.1) with the same initial data $\theta_{1}(0)=\theta_{2}(0)$ and the same control $g$, we have $\theta_{1}=\theta_{2}$ in $\mathbb{X}_{\alpha,\beta}$ as long as $\theta_{1}\in\mathcal{C}_{0}$ and $\theta_{2}$ satisfies the $H^{\alpha-2\beta}$-energy inequality: for every $t\in[0,T],$
@@
[…]
@@
Assume that the law of the initial data $\theta_{\varepsilon,\delta(\varepsilon)}(0)$ is $\mathcal{G}\left(0,\varepsilon Q_{\delta(\varepsilon)}/2\right)$, the Gaussian measure on $\dot{H}^{\alpha-2\beta}(\mathbb{T}^{2})$ with mean zero and covariance $\varepsilon Q_{\delta(\varepsilon)}/2.$
Now we provide the definition of solutions. The following definition can be regarded as a generalization of the Leray solution.
Definition 1.2.
We say that $(\theta_{\varepsilon,\delta(\varepsilon)},W)$, $\left(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\in[0,T]},\mathbb{P}\right)$ is a stochastic generalized Leray solution of (1.3) with initial data $\theta_{0}$ if, $\theta_{0}$ is an $H^{\alpha-2\beta}(\mathbb{T}^{2})$-valued random element, $W$ is a cylindrical Wiener process on $H^{\alpha-2\beta}(\mathbb{T}^{2})$ independent of $\theta_{0}$, and $\theta_{\varepsilon,\delta(\varepsilon)}$ is a progressively measurable process satisfying:
(i) $\mathbb{P}$-almost surely, $\theta_{\varepsilon,\delta(\varepsilon)}\in L^{\infty}([0,T];H^{\alpha-2\beta}(\mathbb{T}^{2}))\cap L^{2}([0,T];H^{2\alpha-2\beta}(\mathbb{T}^{2}))\cap C([0,T];H_{w}^{\alpha-2\beta}(\mathbb{T}^{2}))$.
$\left\langle\theta_{\varepsilon,\delta(\varepsilon)}(t),\varphi(t)\right\rangle+\int_{0}^{t}\left\langle\theta_{\varepsilon,\delta(\varepsilon)}(s),\Lambda^{2\alpha}\varphi(s)\right\rangle\,\mathrm{d}s=\left\langle\theta_{0},\varphi(0)\right\rangle+\int_{0}^{t}\langle\theta_{\varepsilon,\delta(\varepsilon)}(s),\partial_{s}\varphi(s)\rangle\,\mathrm{d}s$
$-\frac{1}{2}\int_{0}^{t}\left\langle R_{2}\theta_{\varepsilon,\delta(\varepsilon)}(s),\left[\Lambda,\partial_{1}\varphi(s)\right]\Lambda^{-1}\theta_{\varepsilon,\delta(\varepsilon)}(s)\right\rangle\,\mathrm{d}s$
$+\frac{1}{2}\int_{0}^{t}\left\langle R_{1}\theta_{\varepsilon,\delta(\varepsilon)}(s),\left[\Lambda,\partial_{2}\varphi(s)\right]\Lambda^{-1}\theta_{\varepsilon,\delta(\varepsilon)}(s)\right\rangle\,\mathrm{d}s+\sqrt{\varepsilon}\int_{0}^{t}\langle\Lambda^{2\beta}\varphi(s),\sqrt{Q_{\delta}(\varepsilon)}\,\mathrm{d}W(s)\rangle.$
Here and subsequently, the nonlinear terms $\left\langle R_{j}\theta_{\varepsilon,\delta(\varepsilon)}(s),\left[\Lambda,\partial_{i}\varphi(s)\right]\Lambda^{-1}\theta_{\varepsilon,\delta(\varepsilon)}(s)\right\rangle(i,j=1,2)$ are defined by the $H^{-1/2}-H^{1/2}$ dual, see Section 2 for details.
(iii) $\theta_{\varepsilon,\delta(\varepsilon)}$ satisfies the pathwise $H^{\alpha-2\beta}$-energy inequality: $\mathbb{P}$-almost surely, for all $t\in[0,T]$,
$\frac{1}{2}\left\|\theta_{\varepsilon,\delta(\varepsilon)}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\left\|\theta_{\varepsilon,\delta(\varepsilon)}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s$ $\leqslant\frac{1}{2}\left\|\theta_{0}\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$ (1.9)
$+\sqrt{\varepsilon}\int_{0}^{t}\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,\delta(\varepsilon)}(s),\sqrt{Q_{\delta(\varepsilon)}}\,\mathrm{d}W(s)\rangle$ $+\frac{\varepsilon}{2}\left\|\Lambda^{\alpha}\circ\sqrt{Q_{\delta(\varepsilon)}}\right\|_{HS}^{2}t.$
@@
For every $\varepsilon,\delta(\varepsilon)>0$, let $\theta_{\varepsilon,\delta(\varepsilon)}$ be a stochastic generalized Leray solution of (1.3) in the sense of Definition 1.2 with initial data $\theta_{\varepsilon,\delta(\varepsilon)}(0)$ satisfying Assumption A. Let $\mu_{\varepsilon}=\mu_{\varepsilon,\delta(\varepsilon)}$ be the laws of $\theta_{\varepsilon,\delta(\varepsilon)}$ on $\mathbb{X}_{\alpha,\beta}$. Let $\mathcal{C}_{0}\subset\mathbb{X}_{\alpha,\beta}$ be a weak-strong uniqueness class of (1.1) and $\mathcal{C}=\overline{\mathcal{C}_{0}}^{\mathcal{I}}$. Assume that the scaling regime (1.7) holds for $(\varepsilon,\delta(\varepsilon))$. Then for any closed set $F\subset\mathbb{X}_{\alpha,\beta}$,
@@
We further comment that the above large deviations result holds for a sequence of Galerkin approximations of (1.3) as well, see Proposition 5.1 and Proposition 6.1 for details. Based on this large deviations result, we provide a direct relationship between regularity classes of the weak-strong uniqueness and regularity classes where the energy equality holds. Concretely, we show that the kinetic energy equality holds on every time-reversible set of a certain closure of the weak-strong uniqueness class $\mathcal{C}_{0}$. This introduces the following result. We introduce the time-reversal operator $\mathfrak{T}_{T}:\mathbb{X}_{\alpha,\alpha/2}\rightarrow\mathbb{X}_{\alpha,\alpha/2}$ defined by $(\mathfrak{T}_{T}\theta)(\cdot):=-\theta(T-\cdot)$.
@@
Assume that $\beta=\alpha/2$. Let $\mathcal{C}_{0}\subset\mathbb{X}_{\alpha,\alpha/2}$ be a weak-strong uniqueness class of (1.1) such that $\mathcal{C}=\overline{\mathcal{C}_{0}}^{\mathcal{I}}$ contains non-empty time-reversible subsets. Let $\mathcal{R}\subset\mathcal{C}$ satisfy $\mathcal{R}=\mathfrak{T}_{T}\mathcal{R}$. Suppose that $\theta\in\mathcal{R}$ is a weak solution of (1.1) for some $g\in L^{2}([0,T];L^{2}(\mathbb{T}^{2}))$ in the sense of Definition 4.1, then the following kinetic energy equality holds:
@@
In the proof of Theorem 1.4, we rely on a time-reversibility argument for solutions of Galerkin approximation equation starting from equilibrium with $\delta=0$ (see Lemma 7.2 below). Hence we consider random initial data rather than deterministic ones. We further remark that in the proof for Theorem 1.3, the Assumption A for initial data could be extended to more general conditions (see [GHW24_LLNS, Assumption 2.1]).
@@
Let $\varepsilon>0$ and $m(\varepsilon)\in\mathbb{N}_{+}$ satisfy the scaling regime (2.5). Let $\theta_{\varepsilon,m(\varepsilon)}$ be the solution of (1.12) with initial data $\theta_{\varepsilon,m(\varepsilon)}(0)\sim\mathcal{G}(0,\varepsilon P_{m(\varepsilon)}/2)$ and let $\mu_{\varepsilon,m(\varepsilon)}$ be the law of $\theta_{\varepsilon,m(\varepsilon)}$ on $\mathbb{X}_{\alpha,\alpha/2}.$ Suppose that $\theta$ is a weak solution of (1.1) for some $g\in L^{2}([0,T];L^{2}(\mathbb{T}^{2}))$ in the sense of Definition 4.1.
@@
$\inf_{\{G\text{ is open in }\mathbb{X}_{\alpha,\alpha/2}:\theta\in G\}}\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon,m(\varepsilon)}(G)<-\mathcal{I}(\theta).$
@@
$\inf_{\{G\text{ is open in }\mathbb{X}_{\alpha,\alpha/2},\,\mathfrak{T}_{T}\theta\in G\}}\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon,m(\varepsilon)}(G)<-\mathcal{I}(\mathfrak{T}_{T}\theta).$
Finally, we provide an analysis on the exact representation of the quasi-potential (1.5). Let $\mathcal{C}_{0}$ be a weak-strong uniqueness class on $0,\infty)$ (See Definition 8.2). Denote
$\mathcal{A}_{0}=\Big\{\theta\in\mathcal{C}_{0}:\partial_{t}\theta\in L^{2}([0,\infty);H^{\alpha-2\beta}(\mathbb{T}^{2})),R^{\perp}\theta\cdot\nabla\theta\in L^{2}([0,\infty);H^{\alpha-2\beta}(\mathbb{T}^{2}))$ ,
$\lim\limits_{t\rightarrow\infty}\|\theta(t)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}=0$ $\Big\}.$